# Quantum Origin of Classical Background Fields from Coherent States

### v1

@@ 8 Conclusion
We have presented a first-principles formulation of QED in which classical background EM fields arise naturally from coherent states of the quantized field. By systematically separating coherent laser modes from other quantum degrees of freedom, our approach unifies operator-based, Furry picture, and path integral descriptions, while maintaining full quantum consistency.
This framework clarifies the conceptual origin of background fields, automatically enforces gauge constraints via the Gupta–Bleuler condition (or equivalently BRST cohomology), and allows treatment of effects beyond the fixed background approximation, including laser depletion and backreaction. Conventional generating functionals and standard strong-field QED formulations appear as well-defined limiting cases within this approach.
The central new elements of this work are the derivation of classical background fields from coherent states at the level of first principles, the construction of QED with coherent background fields within a framework unifying operator-based and path-integral formalisms, and the identification of the fixed background approximation and its extensions, including depletion and backreaction, as natural limiting cases of this construction.
Overall, our results provide a general, intensity-independent foundation for QED in background fields, offering both conceptual clarity and a practical starting point for analyzing high-intensity laser-matter interactions and related strong-field phenomena. In future work, this framework may be extended to study a broader class of quantum processes and to develop systematic approaches for incorporating dynamical background fields in various settings.

### v2

@@ 8 Conclusion
The central result of this work is that QED with prescribed classical background EM fields is not a distinct or modified theory. Rather, it is a boundary-condition limit of full QED, obtained by fixing coherent-state asymptotic boundary conditions of the EM field. This perspective differs from earlier uses of coherent states, in which the background field was treated as an externally prescribed c-number input rather than as a boundary-condition limit of full QED.
Starting from an operator formulation of QED, we have shown how the fixed background field approximation arises naturally as a limiting case of scattering between coherent states, thereby providing a precise and operator-based characterization of its conceptual status within the underlying quantum theory.
Our formulation makes explicit how classical background fields emerge as expectation values of the quantized gauge field in coherent states, while fully preserving the quantum dynamics of the remaining degrees of freedom. Gauge constraints are implemented consistently at the level of the quantized EM field through the Gupta–Bleuler condition, or equivalently through BRST cohomology. Consequently, the coherent-state expectation value defining the classical background field automatically satisfies the corresponding Lorenz-gauge condition. Once identified as a c-number classical field, the background field may therefore be treated in a convenient gauge, without imposing additional physical-state constraints beyond those already enforced in the underlying quantum theory. The time dependence of background fields is shown to arise from the choice of picture, rather than being introduced ad hoc at the level of the Hamiltonian.
Conventional generating functionals and standard formulations of QED with prescribed background fields are recovered as well-defined limiting cases corresponding to fixed coherent-state boundary conditions. Allowing for transitions between different coherent states provides a systematic extension beyond the fixed background approximation, enabling a consistent quantum-mechanical treatment of effects such as depletion and backreaction without double counting of photon degrees of freedom.
The results presented here do not introduce new properties of coherent states themselves, but instead formulate the conventional use of background fields in QED as a controlled and conceptually transparent approximation. The scope of this work is conceptual rather than computational, and we make no claim regarding improvements to perturbative techniques or explicit amplitudes beyond clarifying the operator-level structure underlying QED with prescribed background fields.
This perspective provides a solid foundation for the analysis of strong-field and laser–matter interactions within QED, and for future extensions in which background fields are treated dynamically within a fully quantum framework.